OpenStudy (anonymous):

Simplify. (1/5)(5)(100 + -2 + 2)

OpenStudy (anonymous):

(1/5)(5)(100 + -2 + 2) = 5/5 (100 -2 +2) = 1 ( 100 + 0) = ...?

OpenStudy (anonymous):

well what is the answer

OpenStudy (anonymous):

x..x

OpenStudy (anonymous):

100 + 0 = ?

OpenStudy (anonymous):

ok